# CATBench: A Compiler Autotuning Benchmarking Suite for Black-box Optimization

### v1

@@ 1 Introduction
Bayesian optimization (BO) [35, 29, 45, 18, 20] is a powerful tool for automating the optimization of resource-intensive black-box systems, such as machine learning hyperparameter optimization [16, 42, 28], hardware design [36, 15], and scientific discovery [24, 34, 48]. By intelligently exploring the configuration space and learning from observed performance, Bayesian optimization can efficiently identify high-performing configurations with limited resource consumption.
@@
To facilitate further research and development in Bayesian optimization for autotuning, it is crucial to have a diverse and representative set of benchmarks. A comprehensive benchmarking suite is necessary to evaluate optimization algorithms effectively, ensuring that they can handle the complex and varied nature of real-world autotuning problems. We introduce CATBench [^1] https://github.com/anonymoussisef/catbench, a comprehensive benchmarking suite designed to evaluate black-box optimization algorithms in the context of compiler autotuning. CATBench builds upon the real-world applications used in the BaCO [26] framework, incorporating additional parameters, fidelity levels, and multiple objectives. The suite encompasses a wide range of domains, including tensor algebra, image processing, and machine learning, and targets various backend hardware platforms, e.g., Intel CPUs and Nvidia GPUs. By providing a standardized set of real-world benchmarks with a unified, containerized interface usable on a variety of hardware, CATBench enables fair benchmarking of optimization algorithms and facilitates reproducibility. Additionally, the architecture of our benchmarking suite makes it easy to extend the benchmark suite with new compiler frameworks and autotuning problems.
@@
    Thorough evaluation of popular Bayesian Optimization methods and Evolutionary Algorithms showcasing the properties of the benchmarking suite. Since the existing optimization methods do not cover all the features needed by autotuning we include necessary adaptations for compatibility with the requirements of CATBench.
@@ 2.1 Bayesian Optimization
BO aims to find a minimizer $\bm{x}^{*}\in\mathop{\mathrm{arg\,min}}_{\bm{x}\in\mathcal{S}}f(\bm{x})$ of the black-box function $f(\bm{x}):\mathcal{S}\rightarrow\mathbb{R}$, over some possibly noncontinuous, non-euclidean input space $\mathcal{S}$. Moreover, $f$ can only be observed point-wise and its observations are perturbed by Gaussian noise, $y(\bm{x})=f(\bm{x})+\varepsilon_{i}$, where $\varepsilon_{i}\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2})$.
Gaussian processes (GPs) is the model class of choice in most BO applications. A GP provides a distribution over functions $\hat{f}\sim\mathcal{GP}(m(\cdot),k(\cdot,\cdot))$, fully defined by the mean function $m(\cdot)$ and the covariance function $k(\cdot,\cdot)$. Under this distribution, the value of the function $\hat{f}(\bm{x})$ at a given location $\bm{x}$ is normally distributed with a closed-form solution for the mean $\mu(\bm{x})$ and variance $\sigma^{2}(\bm{x})$. We model a constant mean, so the covariance function fully determines the dynamics $k(\cdot,\cdot)$.
The acquisition function uses the surrogate model to quantify the utility of a point in the search space. Acquisition functions balance exploration and exploitation, typically employing a greedy heuristic. The most common is Expected Improvement (EI) [29, 3], but other acquisition functions, such as Thompson sampling [51] and Upper Confidence Bound [47, 46], are also frequently utilized [17, 55].
@@ 2.3 Related Work
In the context of compiler autotuning, Bayesian Optimization has been successfully applied to find high-performing compiler configurations for various programming languages, compilers, and target architectures [2, 4, 57, 53, 26]. The autotuning problem is formulated as a black-box optimization problem, where the objective function represents the performance metric of interest (e.g., execution time or energy usage) and the input space consists of compiler flags, parameters, and transformations.
@@
In the context of Bayesian optimization, a large collection of benchmarking suites have been proposed. primarily in the context of Hyperparameter Optimization (HPO) [63, 31]. HPOBench [14] provides a collection of models for evaluating hyperparameter optimization algorithms, and captures a range of parameter types as well as multi-fidelity evaluations, whereas [60, 12, 62] target benchmarking of Neural Architecture Search (NAS) algorithms. MCBO [13] collects discrete and categorical tasks and BO algorithms in a modular structure. Lastly, BaCO [26] introduced a set of benchmarks for evaluating Bayesian optimization algorithms in the context of autotuning, derived from real-world applications and compilers, covering a range of domains and target architectures.
@@ 3 The CATBench Benchmarking Suite
CATBench is a comprehensive benchmarking suite designed for evaluating Bayesian optimization algorithms in the context of autotuning. Autotuning presents unique challenges to black-box optimization, including numerous rarely-considered conditions that makes the optimization challenging [26]. CATBench comprises a diverse set of benchmarks derived from real-world applications and compilers, covering various domains, hardware backend platforms, and optimization challenges. The suite is accessible through an intuitive Python interface, enabling users to quickly prototype novel algorithms through the assistance of surrogate models or deploy the benchmarks to a cluster of servers using our client-server architecture.
@@ 3.1 The CATBench Interface
The CATBench Python interface facilitates efficient benchmarking via a client-server architecture and Docker containers. The interface significantly simplifies the evaluation process by providing a consistent environment for performance assessments. The code snippet illustrates a typical benchmarking workflow using CATBench, using either a surrogate model or a hardware benchmarking server:
@@
A naive approach could involve an interface that directly calls the TACO and RISE benchmarks using Python subprocesses. However, this method is inefficient due to the high overhead of repeatedly starting and initializing a benchmark for each configuration, including reloading all input datasets, making the process extremely time-consuming.
To overcome this inefficiency, we designed a client-server benchmark setup. By starting the server once, we can initially load the program and datasets, then continuously listen for and execute new configurations, significantly reducing initialization overhead and streamlining the benchmarking process.
@@
CATBench provides surrogate models of the benchmarks for quick prototyping and piloting. These surrogate models enable preliminary testing and optimization without requiring specific GPUs or CPUs, and can soften particular requirements of the real benchmark. The surrogates aim to provide value in early stages of development, when computational resources may be limited, or quick iterations are needed to inform decisions on algorithm design. We employ Docker containers to encapsulate benchmarking tools and their dependencies, ensuring consistency and reproducibility across diverse environments. This approach effectively enables the user to easily set up an environment that mitigates software version conflicts and system incompatibilities, ensuring reliable and comparable results. This approach, akin to the NASBench work [61] in the context of NAS, is outlined in App. A.
@@ 3.2 Benchmark Characteristics
Table 2: Properties of each benchmark in CATBench. In order, left to right: Exposed parameter types (O=ordinal, C=categorical, P=permutations), number of exposed objectives $M$, input dimensionality $D$, number of fidelity dimensions $F$, number of known valid configurations $|\mathcal{V}\subseteq\mathcal{S}|$, ratio of known valid configurations $\frac{|\mathcal{V}|}{|\mathcal{S}|}$, and the types of constraints included (K=known, H=hidden).
@@ 3.2.1 TACO
The TACO benchmarks focus on sparse matrix, vector, and tensor computations on the CPU. The benchmarks include sparse matrix multiplication (SpMM), sparse matrix-vector multiplication (SpMV), sampled dense-dense matrix multiplication (SDDMM), tensor times vector (TTV), and matricized tensor times Khatri-Rao product (MTTKRP). The search spaces involve ordinal, categorical, and permutation parameters, with known and unknown constraints on how the program can be compiled. The objective is to minimize the execution time and the CPU energy consumption of the generated code. TACO is a compiler written in C++ that generates, compiles, and runs a C program to solve the tensor algebra expression based on the input configuration provided by the user. The optimal choice of values for these optimization parameters will vary based on the specific hardware architecture details of the CPU that the program is running on. See Section B.1 for more details on the TACO benchmarks.
@@ 3.2.2 RISE/ELEVATE
The RISE/ELEVATE benchmarks are derived from the RISE/ELEVATE [49] compiler framework. The RISE programming language and ELEVATE configuration language specify how the framework should generate an optimized GPU OpenCL kernel. The benchmarks cover dense linear algebra (GEMM, Asum, Scal), clustering (Kmeans), and stencil computations (Stencil). The objectives are to minimize execution time of the generated code on the GPU, while minimizing energy on the CPU, the GPU, or both. See Section B.2 for details the benchmark computations. The benchmarks involve ordinal parameters and have a relatively low percentage of valid configurations due to substantial known and hidden constraints. The search spaces involve a mix of ordinal and continuous parameters, with known as well as hidden constraints, which are exposed when a configuration is queried. These benchmarks have moderate search space sizes and multiple fidelity dimensions.
@@ 3.2.3 Multi-fidelity and Multi-objective Benchmarks
Multi-fidelity Evaluation As the evaluations of both energy usage and runtime of compiled programs are typically subject to substantial noise, it is conventional to evaluate a configuration tens of times and use the average (median or mean) observation for optimization purposes. Moreover, there is an additional option to repeat runs several times, where the cache is cleared before re-evaluation to eliminate a source of bias. After evaluation, transitory background noise might appear from periodic background disturbances. To counteract these events, we allow the user to specify the wait time between benchmark repeats. Finally, we also allow for the control of the wait period after the benchmarking is finished before a new configuration may start. This process exposes four fidelity parameters: the number of iterations of the compute the kernel, the re-evaluations after the cache is cleared, waits between repeats, and a wait after evaluation. All benchmarks in CATBench support multi-fidelity evaluation along multiple dimensions, to facilitate additionally complex cost-efficient optimization. The number of fidelity dimensions are displayed in the $F$ column of Table 2.
@@ 4 Results
We now assess various properties of CATBench. We demonstrate the performance of various optimization algorithms on the CATBench, in a single-objective context on TACO, and for multiple objectives on RISE/ELEVATE. We evaluate three optimization algorithms: BaCO, NSGA-II, and Random Search. To provide a diverse display, the TACO benchmarks focus on single-objective optimization for compute time, whereas for the RISE/ELEVATE benchmarks, we perform multi-objective optimization considering both compute time and GPU energy consumption. We assess variability across hardware for three tasks and analyze feature importance for the Stencil and SPMM benchmark. Lastly, we provide insight on multi-objective components of CATBench. Table 3 outlines hardware specifications used for our experiments, which dictates the characteristics of the benchmarks.
@@ 4.2 Understanding the CATBench Benchmarks
Variability Across Hardware We evaluate the variability across hardware for three tasks to assess the general variation, and assess the potential for CATBench in the context of meta-learning. In Figure 4, we plot the empirical PDF for the computational speedup for three tasks: The two TACO tasks SpMV and SDDMM, and the RISE/ELEVATE task Stencil. For the TACO tasks, we run a fixed-seed random search for 5000 iterations on three different types of hardware as specified in Table 3. As displayed in the figure, tuning these application can deliver a 10x improvement in performance from a non-tuned application.
Figure 4: Empirical density of speedups for SPMV, SDDMM and Stencil for three and two sets of hardware, respectively. The distribution of output is similar across tasks, suggesting that a change in hardware yields a similar, albeit not identical task.
Feature Importance Many of the CATBench compilers have similar search spaces, but their characteristics may vary substantially. In Figure 5, the permutation feature importance for both objectives on the Stencil benchmark run on an TitanRTX (RTX) and a TitanV (TV) and the same metric for compute time of the SPMM benchmark across hardware. Some parameters, like the omp_num_threads for SPMM, and the tuned_gs0 and tuned_gs1 for Stencil, have substantially varying feature importance, and the benchmarks are moderately sparse in the number of high-impact dimensions.
@@ 5 Conclusion
We introduce CATBench, a comprehensive benchmarking suite for evaluating Bayesian optimization algorithms in compiler autotuning. CATBench features a diverse set of benchmarks derived from real-world applications, covering various domains, hardware platforms, and optimization challenges. These benchmarks are selected to reflect the unique characteristics of compiler autotuning tasks. Simultaneously, CATBench offers a novel set of challenging tasks to Bayesian optimization through exotic search spaces, multi-objective and multi-fidelity evaluations, and transfer learning capabilities. By providing a standardized set of benchmarks and a unified evaluation interface, CATBench enables fair comparison reproducibility, while enabling accelerated progress towards more efficient and effective autotuning methods.

### v2

@@ 1  Introduction
Bayesian optimization (BO) [37, 31, 47, 18, 20] is a powerful tool for automating the optimization of resource-intensive black-box systems, such as machine learning hyperparameter optimization [16, 44, 28, 29, 30], hardware design [38, 15], and scientific discovery [24, 36, 50]. By intelligently exploring the configuration space and learning from observed performance, BO can efficiently identify high-performing configurations with limited resource consumption.
@@
To facilitate further research and development in BO for autotuning, it is crucial to have a diverse and representative set of benchmarks. A comprehensive benchmarking suite is necessary to evaluate optimization algorithms effectively, ensuring that they can handle the complex and varied nature of real-world autotuning problems. We introduce CATBench [^1] https://github.com/anonymoussisef/catbench, a comprehensive benchmarking suite designed to evaluate black-box optimization algorithms in the context of compiler autotuning. CATBench builds upon the real-world applications used in the BaCO [26] framework, incorporating additional parameters, fidelity levels, and multiple objectives. The suite encompasses a wide range of domains, including tensor algebra, image processing, and machine learning, and targets various backend hardware platforms, e.g., Intel CPUs and Nvidia GPUs. By providing a standardized set of real-world benchmarks with a unified, containerized interface usable on a variety of hardware, CATBench enables fair benchmarking of optimization algorithms and facilitates reproducibility. Additionally, the architecture of our benchmarking suite makes it easy to extend the benchmark suite with new compiler frameworks and autotuning problems.
@@
    Thorough evaluation of popular BO methods and Evolutionary Algorithms showcasing the properties of the benchmarking suite. Since the existing optimization methods do not cover all the features needed by autotuning we include necessary adaptations for compatibility with the requirements of CATBench.
@@ 2.1  Bayesian Optimization
BO aims to find a minimizer $\bm{x}^{*}\in\mathop{\mathrm{arg\,min}}_{\bm{x}\in\mathcal{S}}f(\bm{x})$ of the black-box function $f(\bm{x}):\mathcal{S}\rightarrow\mathbb{R}$, over some possibly discontinuous, non-euclidean input space $\mathcal{S}$. $f$ can only be observed pointwise and its observations are perturbed by Gaussian noise, $y(\bm{x})=f(\bm{x})+\varepsilon_{i}$, where $\varepsilon_{i}\sim\mathcal{N}(0,\sigma_{\varepsilon}^{2})$. The Gaussian processes (GPs) is the model class of choice in most BO applications. A GP provides a distribution over functions $\hat{f}\sim\mathcal{GP}(m(\cdot),k(\cdot,\cdot))$, fully defined by the mean function $m(\cdot)$ and the covariance function $k(\cdot,\cdot)$. Under this distribution, the value of the function $\hat{f}(\bm{x})$ at a given location $\bm{x}$ is normally distributed with a closed-form solution for the mean $\mu(\bm{x})$ and variance $\sigma^{2}(\bm{x})$. We model a constant mean, so the covariance function fully determines the dynamics $k(\cdot,\cdot)$. The acquisition function uses the surrogate model to quantify the utility of a point in $\mathcal{S}$. Acquisition functions balance exploration and exploitation, typically employing a greedy heuristic. The most common is Expected Improvement (EI) [31, 3], but other acquisition functions, such as Thompson sampling [53] and Upper Confidence Bound [49, 48], are also frequently utilized [17, 57].
@@ 2.3  Related Work
In the context of compiler autotuning, BO has been successfully applied to find high-performing compiler configurations for various programming languages, compilers, and target architectures [2, 4, 59, 54, 26]. The autotuning problem is formulated as a black-box optimization problem, where the objective is the performance metric of interest (execution, energy usage) and the input space consists of compiler flags, parameters, and transformations.
@@
In the context of BO, a large collection of benchmarking suites have been proposed. primarily in the context of Hyperparameter Optimization (HPO) [65, 33]. HPOBench [14] provides a collection of models for evaluating hyperparameter optimization algorithms, and captures a range of parameter types as well as multi-fidelity evaluations, whereas [62, 12, 64] target benchmarking of Neural Architecture Search (NAS) algorithms. MCBO [13] collects discrete and categorical tasks and BO algorithms in a modular structure. Lastly, BaCO [26] introduced a set of benchmarks for evaluating BO algorithms in the context of autotuning, derived from real-world applications and compilers, covering a range of domains and target architectures.
@@ 3  The CATBench Benchmarking Suite
CATBench is a comprehensive benchmarking suite designed for evaluating BO algorithms in the context of autotuning. Autotuning presents unique challenges to black-box optimization, including numerous rarely-considered conditions that makes the optimization challenging [26]. CATBench comprises a diverse set of benchmarks derived from real-world applications and compilers, covering various domains, hardware backend platforms, and optimization challenges. The suite is accessible through an intuitive Python interface, enabling users to quickly prototype novel algorithms through the assistance of surrogate models or deploy the benchmarks to a cluster of servers using our client-server architecture.
@@ 3.1  The CATBench Interface
The CATBench Python interface facilitates benchmarking via a client-server architecture and Docker containers. The interface significantly simplifies the evaluation process by providing a consistent environment for performance assessments. The code snippet illustrates a typical benchmarking workflow using CATBench, using either a surrogate model or a hardware benchmarking server:
@@
A naive approach could involve an interface that directly calls the TACO and RISE benchmarks using Python subprocesses. However, this method is inefficient due to the high overhead of repeatedly starting and initializing a benchmark for each configuration, including reloading all input datasets, making the process extremely time-consuming. To overcome this inefficiency, we designed a client-server benchmark setup. By starting the server once, we can initially load the program and datasets, then continuously listen for and execute new configurations, significantly reducing initialization overhead and streamlining the benchmarking process.
@@
CATBench provides surrogate models of the benchmarks for quick prototyping and piloting. These surrogate models enable preliminary testing and optimization without requiring specific GPUs or CPUs, and can soften particular requirements of the real benchmark. The surrogates aim to provide value in early stages of development, when computational resources may be limited, or quick iterations are needed to inform decisions on algorithm design. We employ Docker containers to encapsulate benchmarking tools and their dependencies, ensuring consistency and reproducibility across diverse environments. This approach enables the user to easily set up an environment that mitigates software version conflicts and system incompatibilities, ensuring reliable and comparable results. This approach, akin to NASBench [63] in the context of NAS, is outlined in App. A.
@@ 3.2  Benchmark Characteristics
Table 2: Properties of each benchmark in CATBench. In order, left to right: Exposed parameter types (O=ordinal, C=categorical, P=permutations), number of exposed objectives $M$, input dimensionality $D$, number of fidelity dimensions $F$, number of known valid configurations $|\mathcal{V}\subseteq\mathcal{S}|$, ratio of known valid configurations $\frac{|\mathcal{V}|}{|\mathcal{S}|}$, and the types of constraints (K=known, H=hidden).
@@ 3.2.1  TACO
The TACO benchmarks focus on sparse matrix, vector, and tensor computations on the CPU. The benchmarks include sparse matrix multiplication (SpMM), sparse matrix-vector multiplication (SpMV), sampled dense-dense matrix multiplication (SDDMM), tensor times vector (TTV), and matricized tensor times Khatri-Rao product (MTTKRP). The search spaces involve ordinal, categorical, and permutation parameters, with known and unknown constraints on how the program can be compiled. The objective is to minimize the execution time and the CPU energy consumption of the generated code. TACO is a compiler written in C++ that generates, compiles, and runs a C program to solve the tensor algebra expression based on the input configuration provided by the user. The optimal choice of values for these optimization parameters will vary based on the specific hardware architecture details of the CPU that the program is running on. See subsection B.1 for more details on the TACO benchmarks.
@@ 3.2.2  RISE/ELEVATE
The RISE/ELEVATE benchmarks are derived from the RISE/ELEVATE [51] compiler framework. The RISE programming language and ELEVATE configuration language specify how the framework should generate an optimized GPU OpenCL kernel. The benchmarks cover dense linear algebra (GEMM, Asum, Scal), clustering (Kmeans), and stencil computations (Stencil). The objectives are to minimize execution time of the generated code on the GPU, while minimizing energy on the CPU, the GPU, or both. See subsection B.2 for details the benchmark computations. The benchmarks involve ordinal parameters and have a relatively low percentage of valid configurations due to substantial known and hidden constraints. The search spaces involve a mix of ordinal and continuous parameters, with known as well as hidden constraints, which are exposed when a configuration is queried. These benchmarks have moderate search space sizes and multiple fidelity dimensions.
@@ 3.2.3  Multi-fidelity and Multi-objective Benchmarks
Multi-fidelity Evaluation As the evaluations of both energy usage and runtime of compiled programs are typically subject to substantial noise, it is conventional to evaluate a configuration tens of times and use the average (median or mean) observation for optimization purposes. Moreover, there is an additional option to repeat runs several times, where the cache is cleared before re-evaluation to eliminate a source of bias. After evaluation, transitory background noise might appear from periodic background disturbances. To counteract these events, we allow the user to specify the wait time between benchmark repeats. Finally, we also allow for the control of the wait period after the benchmarking is finished before a new configuration may start. This process exposes four fidelity parameters: the number of iterations to compute the kernel, the re-evaluations after the cache is cleared, wait between repeats, and a wait after evaluation. All benchmarks in CATBench support multi-fidelity evaluation along multiple dimensions, to facilitate complex cost-efficient optimization. The number of fidelity dimensions are displayed in the $F$ column of Table 2.
@@ 4  Results
We now assess various properties of CATBench. We demonstrate the performance of various optimization algorithms on the CATBench, in a single-objective context on TACO, and for multiple objectives on RISE/ELEVATE. We evaluate three optimization algorithms: BaCO, NSGA-II, and Random Search. To provide a diverse display, the TACO benchmarks focus on single-objective optimization for compute time, whereas for the RISE/ELEVATE benchmarks, we perform multi-objective optimization considering both compute time and GPU energy consumption. We assess variability across hardware for three tasks and analyze feature importance for the Stencil and SPMM benchmark. Lastly, we provide insight on multi-objective components of CATBench. Table 3 outlines hardware specifications used for our experiments, which determines benchmark characteristics.
@@ 4.2  Understanding the CATBench Benchmarks
Variability Across Hardware We evaluate the variability across hardware for three tasks to assess the general variation, and assess the potential for CATBench in the context of meta-learning. In Figure 4, we plot the empirical PDF for the computational speedup of 5000 random samples across three tasks (SpMV, SDDMM, and Stencil) and different hardware types, specified in Table 3. Depending on hardware, tuning these application can deliver between a 10x improvement in performance from a non-tuned, default setting application.
Figure 4: Empirical density of speedups for SPMV, SDDMM and Stencil for three anthreed two sets of hardware, respectively. The distribution of output is similar across tasks, suggesting that a change in hardware yields a similar, albeit not identical task.
Feature Importance Many of the CATBench compilers have similar search spaces, but their characteristics may vary substantially. In Figure 5, the permutation feature importance for both objectives on the Stencil benchmark run on an TitanRTX (RTX) and a TitanV (TV) and the same metric for compute time of the SPMM benchmark across hardware. Some parameters, like omp_num_threads for SPMM, and tuned_gs0 and tuned_gs1 for Stencil, have substantially varying feature importance. The benchmarks are moderately sparse in the number of high-impact dimensions.
@@ 5  Conclusion
We introduce CATBench, a benchmarking suite for evaluating BO algorithms in compiler autotuning. CATBench features a diverse set of benchmarks derived from real-world applications, covering various domains, hardware platforms, and optimization challenges. These benchmarks are selected to reflect the unique characteristics of compiler autotuning tasks. CATBench offers a novel set of challenging tasks to BO through exotic search spaces, multi-objective and multi-fidelity evaluations, and transfer learning capabilities. By providing a standardized set of benchmarks and a unified evaluation interface, CATBench enables fair comparison reproducibility, while enabling accelerated progress towards more efficient and effective autotuning methods.